# Slow-phonon control of spin Edelstein effect in Rashba $d$-wave altermagnets

### v1

@@ II.2 Phonon-dressed $d_{x^{2}-y^{2}}$-wave altermagnet
The phonon-dressed quasiparticle dispersion inherits several key features from the underlying electronic structure and the EPC. The first term, $\tilde{\alpha}_{\vec{k}}=\alpha_{\vec{k}}+gQ_{0}$, represents a uniform shift of all bands due to the static lattice distortion $Q_{0}$. The second term, $\tilde{d}_{\vec{k}}=\sqrt{(\beta_{\vec{k}}+\tilde{g}Q_{0})^{2}+\lambda_{\rm R}^{2}k^{2}}$, controls the spin splitting between $s=\pm 1$ bands. Importantly, $\beta_{\vec{k}}+\tilde{g}Q_{0}$ is an even but anisotropic function of momentum: the coupling $\tilde{g}Q_{0}$ modifies its magnitude in a direction-dependent way, preserving inversion symmetry but breaking $C_{4}$ rotational symmetry. As a result, the dressed bands show spin-dependent gaps and anisotropic distortions whose strength grows with $Q_{0}$, explaining the deformation. The Rashba term $\lambda_{\rm R}^{2}k^{2}$ further broadens the spin splitting, ensuring that even when $\beta_{\vec{k}}+\tilde{g}Q_{0}\to 0$ along some directions, the spin bands remain separated by a gap proportional to $|\lambda_{\rm R}k|$.
Since our effective model is valid only within a restricted energy window, we focus on very low temperatures, ensuring that thermal excitations remain within the model’s regime of validity. In the zero-temperature limit, the Fermi-Dirac occupations reduce to $f_{\vec{k},s}\approx\Theta(\mu-\mathcal{E}_{\vec{k},s})$, where $\Theta$ is the Heaviside function. Therefore, $n_{\rm e}$ and $m$ satisfy the self-consistent equations
@@
$m\approx{}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k}}\left[\Theta\left(\mu-\mathcal{E}_{\vec{k},+}(Q_{0})\right)-\Theta\left(\mu-\mathcal{E}_{\vec{k},-}(Q_{0})\right)\right]\,$, (13b)
where $Q_{0}$ depends on both $n_{\rm e}$, and $m$.
@@ III Spin Edelstein effect
In crystals lacking inversion symmetry, spin-orbit interaction allows an external electric field to generate a net nonequilibrium spin or orbital polarization. When it involves electron spin, this effect is referred to as the spin Edelstein effect or Rashba-Edelstein effect [31, 35, 36]. It is important to distinguish between “spin splitting” and “spin polarization” in the context of our study. Spin splitting refers to the intrinsic separation of energy bands for opposite spin projections, which arises from the altermagnetic order and/or RSOC. It is encoded directly in the band structure and exists even in the absence of external perturbations. In contrast, spin polarization denotes an induced magnetization, which arise from electron-phonon interactions in our study. Unlike spin splitting, spin polarization is not a fundamental property of the electronic structure but rather reflects the population imbalance between spin states, which can be strongly modulated by lattice distortions or other interactions. Within the linear response regime, the induced expectation value of a spin operator $\hat{S}^{\ell}$ ($\ell\in\{x,y,z\}$) due to an in-plane electric field $\mathbb{E}^{j}$ along direction $j\in\{x,y\}$ can be written as [37, 38]
@@
where $\chi_{\ell j}$ defines the linear susceptibility, also called the Edelstein response tensor. The response tensor $\chi_{\ell j}$ can be computed using the Kubo formalism. It contains contributions from intraband processes within a single band as well as from interband coherences, given by
$\chi_{\ell j}={}$ $\frac{e}{4\pi^{2}}\int d^{2}k\Bigg[\tau_{\rm intra}\,\sum_{s}\frac{\partial f_{\vec{k},n}}{\partial E_{\vec{k},s}}S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}$
$-i\sum_{s\neq s^{\prime}}\frac{f_{\vec{k},s}-f_{\vec{k},s^{\prime}}}{E_{\vec{k},s}-E_{\vec{k},s^{\prime}}}\frac{S^{\ell}_{\vec{k},s^{\prime}s}\,v^{j}_{\vec{k},ss^{\prime}}}{E_{\vec{k},s^{\prime}}-E_{\vec{k},s}+i/\tau_{\rm inter}}\Bigg]\,.$ (15)
In this study, we set the intraband and interband scattering rates to be equal, $\tau^{-1}_{\rm intra}=\tau^{-1}_{\rm inter}=0.5~\mathrm{eV}$, a value consistent with typical metallic systems [38, 35]. The operator matrix elements are defined as
@@
$v^{j}_{\vec{k},ss^{\prime}}={}$ $\frac{1}{\hbar}\langle\tilde{\psi}_{\vec{k},s}|\partial\mathcal{H}_{\vec{k}}^{\text{SH}}/\partial k_{j}|\tilde{\psi}_{\vec{k},s^{\prime}}\rangle\,.$ (16b)
With the phonon-dressed Hamiltonian from Eq. (9), the corresponding matrix elements can be computed. In the zero-temperature limit, the derivative of the Fermi function reduces to a Dirac delta $\frac{\partial f}{\partial E}\to-\delta(E-\mu)$. Substituting this into the intraband term yields a simplified expression for the low-temperature response $\chi^{\rm intra}_{\ell j}=-\tfrac{e}{4\pi^{2}}\tau_{\rm intra}\int d^{2}k\sum_{s}\delta(E_{\vec{k},s}-\mu)\,S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}$. This form directly connects the intraband Edelstein response to the states at the Fermi surface and their corresponding operator and velocity matrix elements.
In the following, we compute the spin Edelstein effect within our phonon-dressed model, which accounts for the slow lattice dynamics. This allows us to investigate how the phonon-induced renormalization of the chemical potential, $gQ_{0}$, and the generation of an effective magnetization, $\tilde{g}Q_{0}$, influence the spin response. In particular, we analyze how these phonon-mediated modifications compete or cooperate with other intrinsic parameters of the model to shape the magnitude and orientation of the induced spin polarization. In our simulations, for slow phonon, we take the limits $M\to\infty$ and $\omega_{\rm p}\to 0$ such that the product $M\omega_{\rm p}^{2}$, corresponding to the stiffness of the phonon potential, is fixed to unity.
@@ IV Results and discussion
We summarize the key parameters in our model as follows: $g$ (EPC to on-site electron densities), $\tilde{g}$ (EPC to staggered electron densities), $0<\beta<1$ (altermagnetic strength), $\lambda_{\rm R}$ (RSOC strength), and $\mu$ (chemical potential). For the EPCs, we apply the Lindemann criterion [79], which limits atomic displacements to at most 10% of the lattice constant to prevent lattice melting. Specifically, for a representative typical lattice constant of $4-5\,\text{\AA }$ in target $d$-wave altermagnets, we select $g$ such that the equilibrium displacement $Q_{0}$ does not exceed $0.5\,\text{\AA }$. Furthermore, we set $\tilde{g}<g$, since the staggered energy responsible for spin splitting is inherently smaller than the on-site hopping energy. Throughout the subsequent analyses, we set $\tilde{g}=g/4$ unless stated otherwise. Finally, $\lambda_{\rm R}$ and $\mu$ are tuned to cover the bandwidth of the spin channels, ensuring that the spin response is properly aligned with the electronic structure.
In the numerical simulations, we iteratively solve Eqs. (13a) and (13b) for the given EPCs and bare chemical potential $\mu$ [80]. Once $n_{\rm e}$ and $m$ are specified, the equilibrium phonon displacement is given by $Q_{0}=-(g\,[n_{\rm e}-1]+\tilde{g}\,m)/M\omega_{\rm p}^{2}$. Even though $Q_{0}$ appears as a static-Holstein model artifact, it reflects a stable minimum of the full Hamiltonian. It does not decay to zero or diverge, demonstrating that the finite displacement is self-consistent and dynamically robust. In the limit of extremely soft phonon modes ($\omega_{\rm p}\to 0$), the phonon dynamics become much slower than the electronic timescales. Consequently, the kinetic term $P_{i}^{2}/2M$ vanishes, rendering the phonon field effectively static. Our approach can be viewed as a self-consistent, quasi-static treatment of the electron-phonon coupled system, in which the phonon displacement $Q_{0}$ is determined by minimizing the full Hamiltonian for given electronic densities. Unlike a full Holstein model solution, we do not average over multiple phonon configurations or include dynamical phonon fluctuations, and thus many-body effects such as the frequency-dependent self-energy are neglected. While this corresponds formally to the zero-frequency limit of the Migdal perturbation theory at lowest order [81], it is more precise to describe our method as a self-consistent static-Holstein approximation, capturing non-perturbative, density-driven lattice shifts beyond standard linearized treatments.
@@
Once the complete Hamiltonian in Eq. (9) is established, we can apply the Kubo formula from Eq. (III) to calculate the spin Edelstein effect. This approach uses the phonon-dressed Hamiltonian’s eigenvalues and eigenstates to determine the linear response of the spin polarization to an applied electric field.
@@
Figure 2(a) illustrates the evolution of the spin Edelstein susceptibilities, $\chi_{\ell j}/\chi_{0}$, where $\chi_{0}=e/4\pi^{2}$ sets the reference scale, with increasing EPC strength $g$ at $\mu=0$. For weak EPC ($g\lesssim 0.06$ eV/Å), the susceptibilities remains nearly constant. In this regime, electron-phonon scattering is too weak to significantly modify the spin-split band structure, which reflects the perturbative regime. In other words, the spin-momentum locking characteristic of Rashba altermagnet ensures a robust spin Edelstein response in this regime. To explicitly demonstrate the breaking of the antisymmetry relation in the spin Edelstein response, $\chi_{\ell j}\neq-\chi_{j\ell}$, we plot the quantity $\chi_{j\ell}$ for comparison. If the antisymmetric condition $\chi_{\ell j}=-\chi_{j\ell}$ were preserved under the influence of EPC, the curves corresponding to $\chi_{\ell j}$ and $\chi_{j\ell}$ would coincide, exhibiting no deviation. We note that, owing to the symmetry of our model—specifically, $v^{z}_{{\vec{k}},ss^{\prime}}=0$ arising from the absence of a $k_{z}$ component in the Hamiltonian—the spin Edelstein effect cannot generate out-of-plane ($z$-) polarization, leading to $\chi_{\ell z}=\chi_{zj}=0$. Thus, only in-plane spin polarizations are induced.
@@
For all $g$ values, the dominant contributions arise from the off-diagonal elements $\chi_{xy}$ and $\chi_{yx}$. These terms encode the transverse Edelstein response, reflecting the strong spin-momentum locking induced by RSOC. In contrast, the diagonal components $\chi_{xx}$ and $\chi_{yy}$ remain much smaller, with $\chi_{xx}$ slowly decreasing and $\chi_{yy}$ staying nearly constant across a large range of $g$. As $g$ increases, both $\chi_{xy}$ and $\chi_{yx}$ undergo a monotonic decrease, eventually touching zero near $g_{\rm c}\approx 0.25$ eV/Å. This zero response signals the point of spin Edelstein depolarization.
@@
To understand the depolarization phenomenon, we examine how the spin-resolved band structure evolves in the presence of slow static-Holstein phonon. Specifically, we revisit the expression for the spin susceptibility in Eq. (III). The vanishing of $\chi_{\ell j}$ requires the velocity term $v^{j}_{\vec{k}}$ to become zero over the dominant regions of the Brillouin zone. A straightforward scenario arises when the spin-split bands are shifted away from the zero Fermi energy due to $\mu=0$ in our simulations. In such cases, the group velocity $v^{j}_{\vec{k}}=\hbar^{-1}\partial\mathcal{E}_{\vec{k}}/\partial k_{j}$ approaches zero due to the absence of bands near the Fermi energy, rendering the electrons effectively immobile in response to an applied electric field. Consequently, the charge current—and hence the current-induced spin polarization—vanishes.
To identify the critical coupling strength $g_{\rm c}$ at which the spin Edelstein polarization is fully suppressed, marking the onset of depolarization, we consider the band dispersion $\mathcal{E}_{\vec{k},s}$. Since the lower band (s=1s=-1) gives the smallest energies, the condition for the absence of zero-energy states is $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]>0$.
Figure 3 shows the evolution of the minimum of the lower spin-split band as the EPC $g$ is increased. The critical coupling $\tilde{g}_{\rm c}$ emerges when $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]=0$, with its value depending on the specific parameter space. In Fig. 3(a), increasing the altermagnetic anisotropy $\beta$ shifts the curves downward, indicating that stronger anisotropy delays the depolarization transition to higher $\tilde{g}_{\rm c}$. Similarly, in Fig. 3(b), increasing the RSOC $\lambda_{\rm R}$ lowers the minimum band energy, thereby increasing $\tilde{g}_{\rm c}$ and delaying the onset of depolarization. The circles mark the critical coupling $\tilde{g}_{\rm c}$ for each parameter set, defined by the condition $\min[\mathcal{E}_{\vec{k},-}]=\mu$. A slight mismatch arises between $\tilde{g}_{\rm c}$ and the actual onset of depolarization at $g_{\rm c}$ in the Edelstein response. This discrepancy originates from the fact that depolarization sets in only after the zeros of $\min[\mathcal{E}_{\vec{k},-}]$ are crossed, i.e., when the Fermi surface is fully collapsed due to the EPC-induced energy shifts and band anisotropies. However, the EPC contribution determined by the condition $\min[\mathcal{E}_{\vec{k},-}]=\mu$ corresponds to the point where the minimal band energy just reaches the chemical potential, thus, $g_{\rm c}\gtrsim\tilde{g}_{\rm c}$.
Figure 3: Minimum of the lower spin-split band, $\min[\mathcal{E}_{\vec{k},-}]$, as a function of the EPC $g$ for two parameter variations. (a) Variation of the altermagnetic anisotropy $\beta$ and (b) variation of the RSOC strength $\lambda_{\rm R}$. The horizontal dashed line marks the zero Fermi energy $\mu$. Open circles indicate the critical EPC $\tilde{g}_{\rm c}$ at which the lower band touches $\mu$, marking the onset of spin Edelstein depolarization at $g_{\rm c}\gtrsim\tilde{g}_{\rm c}$.
@@
Figure 4 presents the DOS $\mathcal{D}(\mathcal{E})$ along the high-symmetry $\Gamma\to X$ and $\Gamma\to Y$ directions for various values of the EPC $g$. At weak coupling [Fig. 4(a), $g=0.04$ eV/Å], the band dispersions along $X$ and $Y$ are highly symmetric, leading to overlapping isotropic DOS peaks. The sharp peaks in $\mathcal{D}(\mathcal{E})$ correspond to van Hove singularities originating from nearly degenrate states in the band structure. This, in turn, implies that the spin Edelstein susceptibilities satisfy $\chi_{\ell j}=-\chi_{j\ell}$, as illustrated by the spin Edelstein responses in Fig. 2(a) near $g=0.04$ eV/Å, reflecting the intrinsic antisymmetric character of the bands. Upon increasing $g$ to $0.12$ eV/Å  [Fig. 4(b)], a small anisotropy emerges between the $X$ and $Y$ directions, offering slightly antisymmetric breaking feature $\chi_{\ell j}\neq-\chi_{j\ell}$, as shown by the responses in Fig. 2(a) near $g=0.12$ eV/Å.
At $g\approx 0.18$ eV/Å  [Fig. 4(c)], the band dispersions become markedly asymmetric associated with more split peaks. The inset shows that the bands become slightly warped: the spin-down dispersion along $X$ shifts downward relative to $Y$. This anisotropy is directly reflected in the DOS, where a double-peak structure begins to form, indicating the splitting of van Hove singularities associated with anisotropic band edges. The critical EPC, above which permanent depolarization occurs, is $g_{\rm c}\approx 0.25$ eV/Å. For $g=0.28$ eV/Å  [Fig. 4(d)], the DOS peaks are more separated than in the previous case shown in Fig. 4(c). Due to the absence of spin bands at the Fermi energy (see the inset showing the spin-resolved band structure in Fig. 4(d)), the intraband and interband contributions to the spin Edelstein susceptibility in Eq. (III) are strongly suppressed. Thus, electrons are highly localized; they do not propagate easily, which can strongly suppress transport-related properties.

### v2

@@ II.2 Phonon-dressed $d_{x^{2}-y^{2}}$-wave altermagnet
The phonon-dressed quasiparticle dispersion inherits several key features from the underlying electronic structure and the EPC. The first term, $\tilde{\alpha}_{\vec{k}}=\alpha_{\vec{k}}+gQ_{0}$, represents a uniform shift of all bands due to the static lattice distortion $Q_{0}$. The second term, $\tilde{d}_{\vec{k}}=\sqrt{(\beta_{\vec{k}}+\tilde{g}Q_{0})^{2}+\lambda_{\rm R}^{2}k^{2}}$, controls the spin splitting between $s=\pm 1$ bands. Importantly, $\beta_{\vec{k}}+\tilde{g}Q_{0}$ is an even but anisotropic function of momentum: the coupling $\tilde{g}Q_{0}$ modifies its magnitude in a direction-dependent way, preserving inversion symmetry but breaking $C_{4}$ rotational symmetry. As a result, the dressed bands show spin-dependent gaps and anisotropic distortions whose strength grows with $Q_{0}$, explaining the deformation. The Rashba term $\lambda_{\rm R}^{2}k^{2}$ further broadens the spin splitting, ensuring that even when $\beta_{\vec{k}}+\tilde{g}Q_{0}\to 0$ along some directions, the bands remain separated by a gap proportional to $|\lambda_{\rm R}k|$.
Since our effective model is valid only within a restricted energy window, we focus on very low temperatures, ensuring that thermal excitations remain within the model’s regime of validity. In the zero-temperature limit, the Fermi-Dirac occupations reduce to $f_{\vec{k},s}\approx\Theta(\mu-\mathcal{E}_{\vec{k},s})$, where $\Theta$ is the Heaviside function. Therefore, $n_{\rm e}$ and $m_{\rm s}$ satisfy the self-consistent equations
@@
$m_{\rm s}\approx{}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k}}\left[\Theta\left(\mu-\mathcal{E}_{\vec{k},+}(Q_{0})\right)-\Theta\left(\mu-\mathcal{E}_{\vec{k},-}(Q_{0})\right)\right]\,$, (16b)
where $Q_{0}$ depends on both $n_{\rm e}$ and $m_{\rm s}$. Within the static-Holstein approximation, $Q_{0}$ renormalizes the band energies according to Eq. (15), independent of any external electric field. As the EPC increases, the phonon-induced energy shifts can push the spin-split bands entirely above (or below) the chemical potential, leading to a complete disappearance of Fermi-surface states. We refer to this phenomenon as a phonon-induced Fermi-surface collapse. This collapse is an equilibrium effect determined solely by the band renormalization and does not rely on transport or response functions. In the following sections, we show that the vanishing of the Edelstein spin polarization occurs only after this Fermi-surface collapse, demonstrating that the depolarization is a consequence—not the origin—of the equilibrium band reconstruction.
@@ III Spin Edelstein effect
In noncentrosymmetric crystals, spin-orbit interaction enables an external electric field to induce a nonequilibrium spin or orbital polarization. When involving electron spin, this phenomenon is known as the spin Edelstein (Rashba-Edelstein) effect [31, 35, 36]. It is crucial to distinguish between spin splitting and spin polarization. Spin splitting refers to the intrinsic energy separation of bands with opposite spin projections, originating from altermagnetic order and/or RSOC, and is encoded in the equilibrium band structure even in the absence of external perturbations. In contrast, spin polarization denotes an induced magnetization arising from nonequilibrium population imbalance between spin states, generated here by electron-phonon interactions or electric fields. Unlike spin splitting, spin polarization is not an intrinsic property of the electronic structure but is strongly tunable by lattice distortions and external driving.
While Fermi-surface modulations due to the EPC associated with possible metal-to-(band)insulator transitions can be detected using thermodynamic quantities such as the heat capacity, most of these probes are sensitive only to the total density of states. Here, our focus is on how the spin Edelstein response evolves as the Fermi surface shrinks. Unlike conventional thermodynamic measures, the Edelstein tensor provides symmetry-resolved information about the spin structure of the electronic states, making it a particularly sensitive probe of distortion-controlled altermagnetic behavior.
Within the linear response regime, the induced expectation value of a spin operator $\hat{S}^{\ell}$ ($\ell\in\{x,y,z\}$) due to an in-plane electric field $\mathbb{E}^{j}$ along direction $j\in\{x,y\}$ can be written as [37, 38]
@@
where $\chi_{\ell j}$ defines the linear susceptibility, also called the Edelstein response tensor. It contains contributions from intraband processes within a single band as well as from interband coherences, given by
$\chi_{\ell j}={}$ $\frac{e}{4\pi^{2}}\int d^{2}k\Bigg[\tau_{\rm intra}\,\sum_{s}\frac{\partial f_{\vec{k},s}}{\partial\mathcal{E}_{\vec{k},s}}S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}$
$-i\sum_{s\neq s^{\prime}}\frac{f_{\vec{k},s}-f_{\vec{k},s^{\prime}}}{\mathcal{E}_{\vec{k},s}-\mathcal{E}_{\vec{k},s^{\prime}}}\frac{S^{\ell}_{\vec{k},s^{\prime}s}\,v^{j}_{\vec{k},ss^{\prime}}}{\mathcal{E}_{\vec{k},s^{\prime}}-\mathcal{E}_{\vec{k},s}+i/\tau_{\rm inter}}\Bigg]\,.$ (18)
In this study, we set the intraband and interband broadening times to be equal, $\tau^{-1}_{\rm intra}=\tau^{-1}_{\rm inter}=0.5~\mathrm{eV}$, a value consistent with typical metallic systems [38, 35]. We note that all energies appearing in Eq. (III) correspond to the phonon-dressed quasiparticle energies defined in Eq. (15). The operator matrix elements are defined as
@@
$v^{j}_{\vec{k},ss^{\prime}}={}$ $\frac{1}{\hbar}\langle\tilde{\psi}_{\vec{k},s}|(\partial\mathcal{H}_{\vec{k}}^{\text{SH}}/\partial k_{j})|\tilde{\psi}_{\vec{k},s^{\prime}}\rangle\,.$ (19b)
Our evaluation of the Edelstein response employs a constant relaxation-time approximation. In the present static-Holstein framework, the ladder-type vertex corrections vanish within linear response [87, 88]. This is because the phonon coordinate is replaced by its equilibrium value $Q_{0}$, yielding a purely static and quadratic Hamiltonian, with the phonon effects entering only as renormalizations of the single-particle. Thus, the resulting theory does not involve a dynamical phonon propagator, frequency-dependent electron self-energy, or phonon-induced scattering. In this strictly static framework, therefore, there are no phonon-induced vertex corrections in the many-body sense. A full treatment of vertex corrections in the presence of dynamical phonons is beyond our current scope and will be left for future investigation.
@@
With the phonon-dressed Hamiltonian from Eq. (12), the corresponding matrix elements can be computed. In the zero-temperature limit, the derivative of the Fermi function reduces to a Dirac delta $\frac{\partial f}{\partial\mathcal{E}}\to-\delta(\mathcal{E}-\mu)$. Substituting this into the intraband term yields a simplified expression for the low-temperature response
$\chi^{\rm intra}_{\ell j}=-\frac{e}{4\pi^{2}}\tau_{\rm intra}\int d^{2}k\sum_{s}\delta(\mathcal{E}_{\vec{k},s}-\mu)\,S^{\ell}_{\vec{k},ss}\,v^{j}_{\vec{k},ss}\,.$ (20)
This form directly connects the intraband Edelstein response to the states at the Fermi surface and their corresponding operator and velocity matrix elements.
In the following, we compute the spin Edelstein effect within our phonon-dressed model, which accounts for the slow phononic effects. This allows us to investigate how the distortion-induced renormalization of the chemical potential, $gQ_{0}$, and the generation of an effective magnetization, $\tilde{g}Q_{0}$, influence the spin Edelstein polarization. In particular, we analyze how these phonon-mediated modifications compete or cooperate with other intrinsic parameters of the model to shape the magnitude and orientation of the induced spin Edelstein polarization. In our simulations, for slow phonon, we take the limits $M\to\infty$ and $\omega_{\rm p}\to 0$ such that the product $M\omega_{\rm p}^{2}$, corresponding to the stiffness of the phonon potential, is fixed to unity.
Throughout this work, the limit $\omega_{\rm p}\to 0$ should be understood as the adiabatic (static) limit of an internal lattice distortion, rather than as the long-wavelength limit of an acoustic phonon. In other words, the limit $\omega_{\rm p}\to 0$ is taken only in the Born–Oppenheimer sense [89], rather than a dynamical phonon mode softening to zero frequency, ensuring a classical lattice field with fixed stiffness $M\omega_{\rm p}^{2}$. While a uniform acoustic displacement corresponds to a rigid translation and does not affect the electronic spectrum, the phonon mode considered here represents a zone-center optical distortion with relative atomic displacements inside the unit cell. Such a distortion generates a finite local potential and therefore leads to a nonvanishing EPC even in the static limit.
We briefly note that the Edelstein effect, describing current-induced spin polarization via a shift of a spin-textured Fermi surface, remains a useful conceptual framework in our system [33]. Under an applied electric field, the electronic distribution acquires a momentum shift, and the resulting spin polarization arises from the spin expectation values of the occupied Bloch states at the displaced Fermi surface. In the limit of dominant RSOC and negligible altermagnetic splitting, our model smoothly reduces to the conventional Rashba 2D electron gas, reproducing the familiar helical spin texture. In contrast, in the altermagnetic regime with strain-induced symmetry lowering, the spin texture is qualitatively modified. While we do not explicitly show the helical spin configuration in momentum space, through band structure analysis in the following, we illustrate how the momentum-dependent exchange splitting and sublattice symmetry breaking reshape the Fermi surface geometry.
@@ IV Results and discussion
[…]
In the numerical simulations, Eqs. (16a) and (16b) are solved self-consistently for given EPCs and bare chemical potential $\mu$ [91]. For specified $n_{\rm e}$ and $m_{\rm s}$, the equilibrium phonon displacement is $Q_{0}=-\left(g\,[n_{\rm e}-1]+\tilde{g}\,m_{\rm s}\right)/M\omega_{\rm p}^{2}$. Although $Q_{0}$ arises within a static-Holstein framework, it corresponds to a stable minimum of the full Hamiltonian and remains finite and self-consistent. Our approach thus constitutes a self-consistent, quasi-static treatment in which $Q_{0}$ is determined by minimizing the coupled electron-phonon Hamiltonian. In the slow-phonon limit ($\omega_{\rm p}\!\to\!0$), inelastic scattering and dissipative processes are neglected. While dynamical fluctuations and frequency-dependent self-energies are neglected, the method captures nonperturbative, density-driven lattice shifts beyond linearized treatments, corresponding formally to the zero-frequency limit of lowest-order Migdal theory [92]. We stress that including dynamical phonons would introduce self-energy corrections and finite lifetimes, potentially smoothing sharp spectral features, but lies beyond the present scope.
@@
Once the complete Hamiltonian in Eq. (12) is established, we can apply the Kubo formula from Eq. (III) to calculate the spin Edelstein effect. Within this approach, the linear spin response to an applied electric field is computed from the eigenvalues and eigenstates of the phonon-dressed Hamiltonian, thereby probing phonon-induced renormalization effects.
@@
Figure 2(a) illustrates the evolution of the spin Edelstein susceptibilities, $\chi_{\ell j}/\chi_{0}$, where $\chi_{0}=e/4\pi^{2}$ sets the reference scale, with increasing EPC strength $g$ at $\mu=0$. For weak EPC ($g\lesssim 0.06$ eV/Å), the susceptibilities remains nearly constant. In this regime, electron-phonon interaction is too weak to significantly modify the spin-split band structure, which reflects the perturbative regime. In other words, the spin-momentum locking characteristic of Rashba altermagnet ensures a robust spin Edelstein response in this regime. To explicitly demonstrate the breaking of the antisymmetry relation in the spin Edelstein response, $\chi_{\ell j}\neq-\chi_{j\ell}$, we plot the quantity $\chi_{j\ell}$ for comparison. If the antisymmetric condition $\chi_{\ell j}=-\chi_{j\ell}$ were preserved under the influence of EPC, the curves corresponding to $\chi_{\ell j}$ and $\chi_{j\ell}$ would coincide, exhibiting no deviation. We note that, owing to the symmetry of our model—specifically, $v^{z}_{{\vec{k}},ss^{\prime}}=0$ arising from the absence of a $k_{z}$ component in the Hamiltonian—the spin Edelstein effect cannot generate out-of-plane ($z$-) polarization, leading to $\chi_{\ell z}=\chi_{zj}=0$. Thus, only in-plane spin polarizations are induced.
@@
For all $g$ values, the dominant contributions arise from the off-diagonal elements $\chi_{xy}$ and $\chi_{yx}$. These terms encode the transverse Edelstein response, reflecting the strong spin-momentum locking induced by RSOC. In contrast, the diagonal components $\chi_{xx}$ and $\chi_{yy}$ remain much smaller, with $\chi_{xx}$ slowly decreasing and $\chi_{yy}$ staying nearly constant across a large range of $g$. As $g$ increases, both $\chi_{xy}$ and $\chi_{yx}$ undergo a monotonic decrease, eventually touching zero near a threshold coupling $g_{\rm c}\approx 0.25$ eV/Å. This zero response signals the point of spin Edelstein depolarization. The coupling strength $g_{\rm c}$ defines a crossover scale at which the phonon-induced band renormalization removes the Fermi surface. No symmetry change or phase transition is associated with this point. Accordingly, the term “transition” is used here in a pragmatic sense to denote the onset of complete Fermi-surface depletion, rather than a sharp thermodynamic transition. It should be emphasized that the strong suppression of the Fermi surface is not the result of an unrealistically large lattice deformation. As noted before, the Lindemann criterion [90] constrains atomic displacements to $\lesssim 10\%$ of the lattice constant, ensuring that the static lattice distortion remains within a physically reasonable range.
Figure 3: Minimum of the lower spin-split band, $\min[\mathcal{E}_{\vec{k},-}]$, as a function of the EPC $g$ for two parameter variations. (a) Variation of the altermagnetic anisotropy $\beta$ and (b) variation of the RSOC strength $\lambda_{\rm R}$. The horizontal dashed line marks the zero Fermi energy $\mu$. Open circles indicate the threshold EPC $G_{\rm c}$ at which the lower band touches $\mu$, marking the onset of spin Edelstein depolarization at $g_{\rm c}\gtrsim G_{\rm c}$.
@@
To understand the depolarization, we examine the spin-resolved band structure in the presence of slow static-Holstein phonons. Revisiting the spin susceptibility in Eq. (III), vanishing $\chi_{\ell j}$ requires the velocity $v^{j}_{\vec{k}}$ to vanish in the dominant regions of the Brillouin zone. This occurs, for example, when the spin-split bands are shifted away from the Fermi energy due to $\mu=0$ in our simulations. In such cases, the group velocity $v^{j}_{\vec{k}}=\hbar^{-1}\partial\mathcal{E}_{\vec{k}}/\partial k_{j}$ approaches zero because there are no bands near the Fermi level. Consequently, both the charge current and the induced spin polarization vanish.
To identify the threshold coupling strength $g_{\rm c}$ at which the spin Edelstein polarization is fully suppressed, marking the onset of depolarization, we consider the band dispersion $\mathcal{E}_{\vec{k},s}$. Since the lower band (s=1s=-1) gives the smallest energies, the condition for the absence of zero-energy states is $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]>0$.
Figure 3 shows the evolution of the minimum of the lower spin-split band as the EPC $g$ is increased. The threshold coupling $G_{\rm c}$ is defined by $\min_{\vec{k}}[\mathcal{E}_{\vec{k},-}]=0$ and depends on the parameter set. In Fig. 3(a), increasing the altermagnetic anisotropy $\beta$ shifts the curves downward, delaying depolarization to higher $G_{\rm c}$. Similarly, in Fig. 3(b), increasing RSOC $\lambda_{\rm R}$ lowers the minimum band energy, also increasing $G_{\rm c}$. Circles mark $G_{\rm c}$ for each parameter, defined by $\min[\mathcal{E}_{\vec{k},-}]=\mu$. The actual onset of depolarization, $g_{\rm c}$, occurs slightly after $G_{\rm c}$, since depolarization sets in only when the Fermi surface fully collapses due to EPC-induced energy shifts and band anisotropies. Therefore, $g_{\rm c}\gtrsim G_{\rm c}$.
@@
Figure 4 shows the DOS $\mathcal{D}(\mathcal{E})$ along the high-symmetry $\Gamma\to X$ and $\Gamma\to Y$ directions for various EPC values $g$. At weak coupling [Fig. 4(a), $g=0.04$ eV/Å], the band dispersions are nearly symmetric, producing overlapping isotropic DOS peaks corresponding to van Hove singularities from nearly degenerate states. Consequently, the spin Edelstein susceptibilities satisfy $\chi_{\ell j}=-\chi_{j\ell}$, as seen in Fig. 2(a) near $g=0.04$ eV/Å. Increasing $g$ to $0.12$ eV/Å [Fig. 4(b)] introduces slight anisotropy between $X$ and $Y$, weakly breaking the antisymmetry, $\chi_{\ell j}\neq-\chi_{j\ell}$, consistent with the spin Edelstein responses in Fig. 2(a) near $g=0.12$ eV/Å.
At $g\approx 0.18$ eV/Å  [Fig. 4(c)], the band dispersions become markedly asymmetric associated with more split peaks. The inset shows that the bands become slightly warped: the spin-down dispersion along $X$ shifts downward relative to $Y$. This anisotropy is directly reflected in the DOS, where a double-peak structure begins to form, indicating the splitting of van Hove singularities associated with anisotropic band edges. The threshold EPC, above which permanent depolarization occurs, is $g_{\rm c}\approx 0.25$ eV/Å. For $g=0.28$ eV/Å  [Fig. 4(d)], the DOS peaks are more separated than in the previous case shown in Fig. 4(c). Due to the absence of bands at the Fermi energy (see the inset showing the spin-resolved band structure in Fig. 4(d)), the intraband and interband contributions to the spin Edelstein susceptibility in Eq. (III) are strongly suppressed.